QUESTION 3 3 p A health club has 4 employees who work on lead generation. Each employee works 35 hours a week, and is paid $17 an hour. Each employee identifies an average of 205 possible leads a week from a list of 8,000 names. Approximately 10 percent of the leads become members and pay a one-time fee of $120. Material costs are $206 per week, and overhead costs are $1,055 per week. Calculate the multifactor productivity for this operation in fees generated per dollar of input. The multifactor productivity for this operation in fees geterated per dollar of nput (round to 4 decimal places)
Answer #1

Multifactor productivity formula = units of output units of labour + units of capital +units of material

units of output = 10% of leads = 205*10% = 20.5 per employee = 20.5*4 = 82 members

members will pay $120 as one time fee total units of output = 82*120 = $9840

units of labour = 4*$17*35 = $2380

units of capital = 206+1055 = $1261

Multifactor productivity = 9840 23801261 = 2.7026 fees per dollar of input.